OpenStudy (anonymous):

Find the slope of the graph of the equation. y= -3x - 7

OpenStudy (anonymous):

@nirmalnema

OpenStudy (anonymous):

@happytales

OpenStudy (anonymous):

This equation can be written as y = (-3)x + (-7). It is a lot like y=mx + c, where m is the slope. So, (-3) should be the slope

OpenStudy (anonymous):

The slope is -3. y=mx+b m is the slope

OpenStudy (anonymous):

so -3 is the answer ?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

thanks